Two equal charges are separated by 3.7 x 10^-10 m. The force between the charges has a magnitude of 2.37 x 10^-3 N. What is the magnitude of q on the charges?

can some one explain to me how to do this step by step. Not that great at math hope you can help. Thanks
Answer #1
q = R sqrt(F/k)
= 3.7*10^-10m*sqrt(2.636*10^-13)C/m
C stands for Coulombs
I typed M where it should be C
answered by: Blish Jackson
Answer #2
Thanks!
answered by: sangeetha
Answer #3
There is a formula for the force (F) between two charges (Q1 and Q2), separated by a distance R. It is called Coulomb's Law. It must be in your text or notes; otherwise they would not have assigned this problem.

The formula says

F = k Q1*Q2/R^2

where k is a constant that equals 8.99*10^9 N*m^2/C^2

In your case, Q1 = Q2 = q, so

F = k*q^2/R^2

Rearrange the equation algebraically and solve for q

q = R sqrt(F/k)= 3.7*10^-10m*sqrt(2.636*10^-13)M/m
= 1.9*10^-16 Coulombs
answered by: Leikesha